# UniMS-RAG: A Unified Multi-source Retrieval-Augmented Generation for Personalized Dialogue Systems

### v1

@@ 7.3.3. Compared with simple planning strategies
 + Ground Retrieval 18.76 20.86 65.89 37.30 41.71 81.53 66.64
 + Ground P & R 19.44 21.89 71.29 37.93 42.67 95.52 87.63
w/o Attention Mask 18.32 20.20 62.70 28.69 31.44 74.53 40.93
w/o Planning
@@
 always do not use source 17.40 19.20 69.78 26.99 28.91 14.73 24.89
7.4. Human Evaluation
Human evaluation is conducted to evaluate the quality of generated response in terms of three metrics: coherence score (Coh.), persona consistency score (Per.Cs for both DuLeMon and KBP), and knowledge consistency score (Know.Cs only for KBP)  [^10] It is worth noting that Per.Cs and Know.Cs here only stand for the first case in Eq. 11.. We randomly sample 100 responses from each dataset with grounding information for each model (We choose UniMS-RAG w/ DPR or w/ ChatGPT and using itself as retriever) and ask three well-educated annotators [^11] The human inspection and annotation were conducted by a reputable data annotation company, and the annotators are compensated fairly based on the market price without revealing any personal information to indicate its coherence score (1-5) and whether the response is consistent with the given persona (1/0), and knowledge (1/0). Table 8 shows the final result. When the number of retrieved evidence is one, we observe there is a significant improvement on DuLeMon and a slight improvement on KBP datasets when using UniMS-RAG w/ DPR. Moreover, when the number of retrieved evidence increases, it is obvious that the performance is largely improved. This observation reveals the effectiveness and robustness of UniMS-RAG to denoise unrelated information, and capture relevant evidences, thanks to the introduction of task and unique training strategies. In addition, UniMS-RAG w/ ChatGPT performs slightly worse than UniMS-RAG w/ DPR due to worse similarity signals. We also observe that human is more likely to find persona-inconsistent cases (Wang et al., 2023a). There are some responses that have intra-sentence or intra-context contradictions (Zheng et al., 2022), for example, the dialogue responds: ”Yes, our license plate starts with ‘Yun A’” to the last user turn: ”Does your license plate start with ‘Yun B’?”.
Table 8. The results of human evaluation. The inter-agreement is about 82%.
Model DuLeMon KBP
Coh. Per.Cs (%) Coh. Per.Cs (%) Know.Cs (%)
FoCus - - 3.53 54.2 33.8
SAFARI 2.84 14.0 4.06 68.0 59.1
UniMS-RAG Number of retrieved evidence = 1
[…]
 w/ ChatGPT 3.35 23.0 4.10 65.2 54.3
Number of retrieved evidence = 3
[…]
 w/ ChatGPT 3.46 61.3 4.15 83.2 61.4
7.5. Case Study
Figure 5. The case study (No cherry-picking)
@@ 7.5. Case Study
7.6. Error Analysis and Limitations
@@
  • Error Propagation. This is a typical issue of solving multiple tasks in sequential order. As shown in Table 7, there is still a gap (i.e., 15% - 20%) between the original UniMS-RAG and the one providing ground-truth planning and retrieval. To avoid this issue, there are some techniques studied currently, including contrastive objectives (Peng et al., 2021) and noisy channel models (Liu et al., 2021).
@@
  • Better Planning and Retrieval. As shown in Table 5 and analysis, better planning and retrieval always lead to better performance, such as the increased number of retrieved evidences, and additional self-refinement mechanism. We also successfully demonstrate that LLMs themselves can be used as retriever simultaneously. Since we observe UniMS-RAG performs dramatically with different dialogue data and different similarity signals, several possible directions are to study how to get more large scale high-quality data or investigate different interactions (i.e, more complex prompting) (Shao et al., 2023; Wang et al., 2023e) to further enhance the performance of planning and retrieval.
@@
  • The Organization of Multiple Sources. We mainly target on four sources of knowledge in this paper (i.e, User-Per, Bot-Per, DOCUMENTS, NULL), however, there are many other useful sources in practice (Wang et al., 2023c), such as user memory (Zhong et al., 2023), local database (Wang et al., 2020) and so on. In addition, the organization of different sources of knowledge also plays a key role in the RAG, such as the different data format (Wu et al., 2021; Ni et al., 2023), or different relationships within sources (Wang et al., 2023e).
8. Conclusion
@@
In this paper, we focus on personalized knowledge-grounded dialogue tasks in a multi-source setting and decompose the problem into three sub-tasks: knowledge source selection, knowledge retrieval and response generation. We discern a notable gap in none of the existing literature concerning the multiple sources planning and auto-regressive retriever with LLMs themselves as backbone. To fill this gap, we propose a Unified Multi-Source Retrieval-Augmented Dialogue System (UniMS-RAG), aiming to build a unified personalized dialogue system with LLM serving as planner, retriever and reader simultaneously. Specifically, we introduce acting tokens to determine which source to call, and evaluation tokens to evaluate the similarity between dialogue context and current retrieved evidence. We explore different similarity signals from DPR or LLMs with evidence mask mechanism to empower LLMs to distinguish the differences between different evidences and filter noises during the training. In addition, we also introduce self-refinement during the inference to iteratively refine the generated responses by leveraging the consistency scores and similarity scores. Experimental results on two popular personalized datasets show that the UniMS-RAG framework can generate more personalized and factual responses and establish a better performance with self-refinement during inference, significantly outperforming strong baseline models under both automatic and human evaluations.
This work is the first attempt toward a unified multi-source retrieval-augmented generation framework for PerDS. There are several limitations and room for further improvement. As discussed in Section 7.6, future works can focus on 1) addressing the error propagation of individual steps; 2) further improving the performance of planning and retrieval; and 3) investigating the organization of different sources of knowledge and incorporating more knowledge (Wang et al., 2023c). We will continue to explore better architectures that leverage large language models for these challenges in personalized dialogue tasks.

### v2

@@ 7.3 Ablation Study
Upper bound of UniMS-RAG. First, we note that the inclusion of ground-truth planning labels or retrieval results casts a positive impact on performance. Planning primarily enhances consistency scores, while grounding retrieval results contributes to BLEU1 and Rouge-L scores. The best results are obtained when both signals are combined, serving as the upper bound of our proposed UniMS-RAG.
The effects of evidence attention mask. There are two notable advantages of adding the evidence attention mask: 1) removing the noise of irrelevant evidence to predict the similarity score; and 2) the UniMS-RAG can effectively operate as a retriever, as it directly captures the relationship between individual contexts and corresponding evidence, optimizing its retrieval capabilities. We can clearly find that removing the evidence attention mask leads to performance degradation, especially on KBP dataset since there are more cases that require external sources of knowledge.
Compared with simple planning strategies. We additionally conduct experiments by adopting two simple planning strategies: 1) always use all sources, User-Per and Bot-Per in DuLeMon, and Both in KBP; and 2) always do not use sources, to investigate the effectiveness of the introduced planning step. The results show that the performance is decreased by 63.18%-1.23% on DuLeMon and 42.07%-36.21% on KBP respectively when using all sources. This highlights that indiscriminate use of all sources not only significantly hampers performance when the majority of samples do not necessitate external references (DuLeMon), but also results in a decline even when a substantial portion of samples requires the utilization of sources (KBP). Conversely, a parallel trend is observed when abstaining from using any external sources. When the majority of samples do not require external references, refraining from their usage leads to notable improvement. However, this approach adversely impacts performance when external sources are essential.
7.4 Human Evaluation
Human evaluation is conducted to evaluate the quality of generated response in terms of three metrics: coherence score (Coh.), persona consistency score (Per.Cs for both DuLeMon and KBP), and knowledge consistency score (Know.Cs only for KBP). We randomly sample 100 responses from each dataset with grounding information for each model (We choose UniMS-RAG w/ DPR or w/ ChatGPT and using itself as retriever) and ask three well-educated annotators [^6] The human inspection and annotation were conducted by a reputable data annotation company, and the annotators are compensated fairly based on the market price without revealing any personal information to indicate its coherence score (1-5) and whether the response is consistent with the given persona (1/0), and knowledge (1/0). Table VIII shows the final result. When the number of retrieved evidence is one, we observe there is a significant improvement on DuLeMon and a slight improvement on KBP datasets when using UniMS-RAG w/ DPR. Moreover, when the number of retrieved evidence increases, it is obvious that the performance is largely improved. This observation reveals the effectiveness and robustness of UniMS-RAG to denoise unrelated information, and capture relevant evidences, thanks to the introduction of task and unique training strategies. In addition, UniMS-RAG w/ ChatGPT performs slightly worse than UniMS-RAG w/ DPR due to worse similarity signals. We also observe that human is more likely to find persona-inconsistent cases [11]. There are some responses that have intra-sentence or intra-context contradictions [58], for example, the dialogue responds: ”Yes, our license plate starts with ‘Yun A’” to the last user turn: ”Does your license plate start with ‘Yun B’?”.
TABLE VIII: The results of human evaluation. The inter-agreement is about 82%.
@@ 7.4 Human Evaluation
7.5 Case Study
Figure 6: The case study (No cherry-picking). We use orange to indicate acting tokens and blue to indicate evaluation tokens.
@@ 7.5 Case Study
7.6 Error Analysis and Limitations
@@
  • Error Propagation. This is a typical issue of solving multiple tasks in sequential order. As shown in Table VII, there is still a gap (i.e., 15% - 20%) between the original UniMS-RAG and the one providing ground-truth planning and retrieval. To avoid this issue, there are some techniques studied currently, including contrastive objectives [59] and noisy channel models [60].
@@
  • Better Planning and Retrieval. As shown in Table V and analysis, better planning and retrieval always lead to better performance, such as the increased number of retrieved evidences, and additional self-refinement mechanism. We also successfully demonstrate that LLMs themselves can be used as retriever simultaneously. Since we observe UniMS-RAG performs dramatically with different dialogue data and different similarity signals, several possible directions are to study how to get more large scale high-quality data or investigate different interactions (i.e, more complex prompting) [61, 35] to further enhance the performance of planning and retrieval.
@@
  • The Number and Organization of Different Sources. We mainly target on four sources of knowledge in this paper (i.e, User-Per, Bot-Per, DOCUMENTS, NULL), however, there are many other useful sources in practice [30], such as user memory [62], local database [63] and so on. In addition, the organization of different sources of knowledge also plays a key role in the RAG, such as the different data format [25, 64], or different relationships within sources [35].
8 Conclusion
@@
In this paper, we focus on personalized knowledge-grounded dialogue tasks in a multi-source setting and decompose the problem into three sub-tasks: knowledge source selection, knowledge retrieval and response generation. We discern a notable gap in none of the existing literature concerning the multiple sources planning and auto-regressive retriever with LLMs themselves as backbone. To fill this gap, we propose a Unified Multi-Source Retrieval-Augmented Dialogue System (UniMS-RAG), aiming to build a unified personalized dialogue system with LLM serving as planner, retriever and reader simultaneously. Experimental results on two popular personalized datasets show that the UniMS-RAG framework can generate more personalized and factual responses and establish a better performance with self-refinement during inference, significantly outperforming strong baseline models under both automatic and human evaluations.
Acknowledgments
This paper was partially supported by grants from the RGC General Research Funding Scheme (GRF) 14222922 (CUHK 2151185). Work done when Hongru Wang is visiting The University of Edinburgh.